# iTACO: Interactable Digital Twins of Articulated Objects from Casually Captured RGBD Videos

### v2

@@ 4.3 Geometry Estimation Results
We evaluate the geometry of our outputs using the Chamfer Distance to the ground truth. We sample 10000 points from the reconstructed mesh and the ground truth mesh, and compute the Chamfer Distance between them. Tab.˜2 shows the results onon S-Dataset. Results on L-Dataset are in the supplement. Due to retrieval errors, link placement errors, and affordance detection errors, Articulate Anything has a high reconstruction error, particularly on movable parts. RSRD performs even worse on since it struggles with clustering 3D Gaussian splats into different parts.
@@ 4.5 Real-World Qualitative Evaluation
We also evaluate our methods and baselines on real data. We use an iPhone 12 Pro with LiDAR camera to capture data. We use Polycam [37] to reconstruct the object surface and Record3D [40] to record the RGBD video of the human interaction with the object. To improve depth map quality, we use Prompt Depth Anything [22] to scale up the original depth map. We compute the image feature matching between the first frame of the video and all the images of the reconstruction stage with LoFTr [46]. We select the image pair with the most reliable feature matching, and compute the rigid transformation between corresponding points of the image pair to align the first video frame to the object surface coordinate system. Real videos include human hands, which interfere with MonST3R predictions for the moving map. Thus, we use Grounded SAM 2 [41, 42] to mask hands with the text prompt: “hands and arms”.
From the qualitative results shown in Fig.˜5, we see that Articulate Anything fails on the Book test case due to the absence of similar meshes in the mesh library. In the Cabinet test case, Articulate Anything retrieves reasonable meshes but predicts the wrong joint rotation direction, as the door rotates into the cabinet. RSRD fails to produce valid results due to the clustering and tracking issues on 3D gaussian. This demonstrates our method can work on real data and outperform current methods in the challenging setting of casually captured real-world RGBD video inputs. See the supplement for more results.
@@ 5 Conclusion
We introduced a new problem setting for reconstructing articulated objects from casually captured RGBD videos. Our problem setting is more practical than prior works, bringing new challenges to current methods. We then developed iTACO: a coarse-to-fine framework for this problem. To evaluate our method, we collected a new dataset that contains 20 more objects than previous benchmarks. Our experiments show our method outperforms baselines, providing a practical and robust approach to acquiring digital twins of articulated objects. Nonetheless, there is still space for improvement. Better dynamic scene understanding approaches and video segmentation models can help improve our method. In addition, our pipeline does not reconstruct the interior of the object. Incrementally reconstructing the interior of the object is an interesting direction for future work.
Acknowledgement
This work was funded in part by a Canada Research Chair, NSERC Discovery Grant, and enabled by support from the Digital Research Alliance of Canada. The authors would like to thank Jiayi Liu, Xingguang Yan, Austin T. Wang, Hou In Ivan Tam, Morteza Badali for valuable discussions, and Yi Shi for proofreading.
@@ A.1 Method
Since the input video usually contains more than 100 frames, we sample the input video to around 20 frames. We first send these sampled frames to MonST3R [82] to compute the moving map of each frame. Then, we select pairs of frames to compute the feature matching using LoFTr [46]. Not all feature matches are accurate, so we only consider feature matches that have a confidence value larger than 0.95 to be valid. We first use feature matching within the static regions to compute the relative camera poses from frame to sequentially. We use these relative camera poses to align all the frames to the first camera coordinate.
Then, we proceed to estimate joint parameters. We select frame pairs within 3 time steps to ensure the accuracy of feature matching. Due to the inaccuracy of moving map and feature matching, we filter out frame pairs that contain fewer than 80 valid feature matches. For the rest of the frame pairs, we estimate joint parameters for the revolute joint and the prismatic joint separately. For each joint type, we run RANSAC [7] for 50 iterations to select the joint parameters estimation that best fits the observation. Based on the estimated joint parameters, we can transform the feature points to expect position at time step . By computing the distance between the expected position and the observed position , we can evaluate how well this set of joint parameters describes the observation. If the distance for the revolute joint is lower, this frame pair votes for revolute, and vice versa. After estimating joint parameters for all the frame pairs, we compute the mean value for the joint axis and position for both joint types. We choose the joint type that gets the most votes as the predicted joint type for coarse prediction.
@@
Though we predict the joint type in coarse prediction, we still retain the prediction results of both joint types from coarse prediction and initialize the refinement module with those values separately. For moving vector initialization, we compute the intersection area between each part and the moving map from MonST3R [82], and then normalize it by the area of each part. We run the forward computation described in the main paper. But computing the Chamfer Distance will be very slow if the number of points is very high. Therefore, for each frame, we randomly sample 50% of points from the whole observation to compute the Chamfer Distance. The gradient will not be affected since we are optimizing the moving vector instead of the probability of being a moving part of each pixel at each frame. This strategy significantly improves the optimization speed without losing prediction accuracy. We concurrently optimize the joint prediction for both joint types for 400 iterations, using Adam optimizer [15] and 5e-3 learning rate.
After optimization, we have two sets of optimized joint parameters, and we need to choose the joint type. We first classify the test case as a prismatic joint if the distance between the object surface point cloud and the joint axis is larger than 0.1 meters. For the remaining cases, we compare the lowest chamfer distance of these two joint types and select the joint type that has the lower Chamfer Distance. These are the final joint type prediction results for our method. Finally, we find pixels in the moving map whose value is larger than 0.7 to be the moving points. We project these points back to 3D and find points in the object surface point cloud whose distance to the moving points is less than 0.03 meters to be the movable parts of the object. The other points in the object surface point cloud are classified as static parts of the object. We reconstruct meshes using NKSR [9].
@@ A.2 Baselines
Robot-See-Robot-Do.
We follow the original pipeline of RSRD to run the experiments on our dataset. RSRD will estimate the motion of each part at each frame. Based on the part motion, we select the motion of the part that has the largest translation distance throughout the video to compute the joint parameters. For geometry reconstruction, we export the gaussian splats from the dig model in RSRD pipeline [14] and build mesh using NKSR [9].
Articulate Anything.
We follow the original pipeline of Articulate Anything to run the experiments on our dataset. Articulate Anything will retrieve top-k objects from the Partnet-Mobility dataset based on the input video. Then, it selects the best one from top-k candidates [18]. When evaluating its results, we consider them to be a failure if the correct object does not lie in the retrieved top-k candidates.
@@ Appendix B Dataset Details
We build our dataset in the SAPIEN simulator [67]. We design a small cuboid space with a simple texture as the environment. Then, we place the object in that space and render videos while manipulating the object using the built-in controller. Fig.˜B.1 illustrates the data collection environment. The statistics are listed in Tab.˜B.1.
| Category | #Object | Joint Type | #Video(revolute, prismatic) | # Average Frames |
@@ C.1 Joint Estimation and Geometry Reconstruction Results on L-Dataset
We conduct a more extensive evaluation of methods, excluding RSRD, on the L-Dataset. Quantitative results on joint estimation and geometry reconstruction are shown in Tab.˜C.2 and Tab.˜C.3. Similar to the results in S-Dataset, our method outperforms Articulate Anything and other ablated versions consistently on L-Dataset.
@@ C.2 Variance of Joint Estimation Results
| Joint | Methods | Axis(rad) | Position(m) | Type | State(rad or m) | Failure |
@@ C.3 Moving Map Segmentation Results
Moving map segmentation is an important intermediate results of our method. We evaluate moving map segmentation using mIOU metric. From the quantitative results in Tab.˜C.5, we find our method achieves comparable results to MonST3R [82], though our pipeline ignores the inner parts of the object, which could be moving part in the video. Without coarse prediction as initialization, our method is hard to optimize for very accurate moving map segmentation. Similarly, without automatic mask generation, our pipeline cannot optimize moving map of the video efficiently, resulting in worse moving map segmentation of the video.
@@ C.4 Camera Pose Estimation Results
There are already several dynamic scene reconstruction works that can infer camera poses of a video sequence. This comparison aims to demonstrate why not directly use camera estimation results from existing dynamic scene reconstruction models. We compare our camera pose prediction results with two very recent works MonST3R [82] and CUT3R [57]. We transform all the camera poses to the camera coordinate of the first video frame. We compare both camera rotation error and translation error on our dataset. From the results shown in Tab.˜C.6, we can find that our method can produce more accurate camera pose estimation, particularly camera orientation, against recent works on dynamic scene understanding. We further evaluate how camera pose estimation will affect the final joint parameter estimation results on S-Dataset. In Tab.˜C.7, we find that using camera poses estimated by MonST3R will lead to a large degradation in the performance.
@@ C.5 Moving Map Prediction Choice
In our coarse prediction module, we use MonST3R [82] to generate a moving map of the input video. A natural question is, why not use an off-the-shelf motion segmentation method? Here we compare the output of moving map prediction results from MonST3R and Segment Any Motion [10], which is a very recent work on motion segmentation on videos. In Fig.˜C.2 we can find that Segment Any Motion produces much larger error prediction results compared to MonST3R. Segment Any Motion first computes point tracks of the video, and selects moving points based on point tracks. Then, it uses these points to prompt Segment Anything 2 [39] to generate the moving map of the video [10]. But in this case, the moving points lie around the edge of the door of this dishwasher. Thus, some points prompt selecting the dishwasher’s top as the moving part instead of the front door. This small position error will lead to a large error in moving map segmentation. In contrast, MonST3R computes the moving map based on optical flow and estimated camera poses. It subtracts the optical flow caused by camera motion from the final optical flow to determine which part of the video is moving. This strategy turns out to be much robust to point-track based motion segmentation methods.
@@ C.6 Discussion on Automatic Video Segmentation Strategy
Besides, we test our method using ground truth mesh-level segmentation on the input video on S-Dataset. An image is rendered from an object comprising several meshes. The mesh-level segmentation identifies different regions that are rendered from different meshes. The results in Tab.˜C.8 show that at least using automatic part segmentation is comparable to using ground truth mesh-level segmentation.
@@ C.7 Additional Real-World Results
In the main paper, we demonstrate the results of building interactable digital twins of real articulated objects with revolute joints. Here we demonstrate reconstructing articulated objects with prismatic joints in Fig.˜C.4. Articulate Anything still struggles with retrieving the correct object in the Storage case, while RSRD fails to crop and cluster the object from the environment for the Drawer case.
C.8 Failure Case Analysis
@@
From the failure modes discussed above, we believe geometry information will help to refine the segmentation results. For example, we can use recent 3D segmentation approaches like SAMPart3D [79] and PartField [26] to produce segmentation on the observed point cloud. This geometry-based segmentation can correct the errors in automatic part segmentation, which is purely image-based.

### v3

@@ 4.3 Geometry Estimation Results
We evaluate the geometry of our outputs using the Chamfer Distance to the ground truth. We sample 10000 points from the reconstructed mesh and the ground truth mesh, and compute the Chamfer Distance between them. Tab.˜2 shows the results on S-Dataset. Results on L-Dataset are in the supplement. Due to retrieval errors, link placement errors, and affordance detection errors, AA has a high reconstruction error, particularly on movable parts. RSRD performs even worse, struggling with clustering 3D Gaussian splats into different parts.
@@ 4.5 Real-World Qualitative Evaluation
We also evaluate our methods and baselines on real data. We use an iPhone 12 Pro with LiDAR camera to capture data. We use Polycam [polycam] to reconstruct the object surface and Record3D [record3d] to record the RGBD video of the human interaction with the object. To improve depth map quality, we use Prompt Depth Anything [lin2025prompting] to scale up the original depth map. We compute the image feature matching between the first frame of the video and all the images of the reconstruction stage with LoFTr [sun2021loftr]. We select the image pair with the most reliable feature matching, and compute the rigid transformation between corresponding points of the image pair to align the first video frame to the object surface coordinate system. Real videos include human hands, which interfere with MonST3R predictions for the moving map. Thus, we use Grounded SAM 2 [GroundedSAM2, ren2024grounded] to mask hands with the text prompt: “hands and arms”.
From the qualitative results shown in Fig.˜5, we see that AA fails on the Book test case due to the absence of similar meshes in the mesh library. In the Cabinet test case, AA retrieves reasonable meshes but predicts the wrong joint rotation direction, as the door rotates into the cabinet. RSRD fails to produce valid results due to the clustering and tracking issues on 3D gaussian. This demonstrates our method can work on real data and outperform current methods in the challenging setting of casually captured real-world RGBD video inputs.
@@ 5 Conclusion
We introduced a new problem setting for reconstructing articulated objects from casually captured RGBD videos. Our problem setting is more practical than prior works, bringing new challenges to current methods. We then developed iTACO: a coarse-to-fine framework for this problem. To evaluate our method, we collected a new dataset that contains 20 more objects than previous benchmarks. Our experiments show our method outperforms baselines, providing a practical and robust approach to acquiring digital twins of articulated objects. Nonetheless, there is still space for improvement. Better dynamic scene understanding approaches and video segmentation models can help improve our method. In addition, our pipeline does not reconstruct the interior of the object. Reconstructing the interior of the object is an interesting direction for future work.
Acknowledgements. This work was funded in part by a Canada Research Chair, NSERC Discovery Grant, and enabled by support from the Digital Research Alliance of Canada. The authors would like to thank Jiayi Liu, Xingguang Yan, Austin T. Wang, Hou In Ivan Tam, Morteza Badali for valuable discussions, and Yi Shi for proofreading.
@@ A.1 Method
Since the input video usually contains more than 100 frames, we sample the input video to around 20 frames. We first send these sampled frames to MonST3R [zhang2024monst3r] to compute the moving map of each frame. Then, we select pairs of frames to compute the feature matching using LoFTr [sun2021loftr]. Not all feature matches are accurate, so we only consider feature matches that have a confidence value larger than 0.95 to be valid. We first use feature matching within the static regions to compute the relative camera poses from frame to sequentially. We use these relative camera poses to align all the frames to the first camera coordinate.
Then, we proceed to estimate joint parameters. We select frame pairs within 3 time steps to ensure the accuracy of feature matching. Due to the inaccuracy of moving map and feature matching, we filter out frame pairs that contain fewer than 80 valid feature matches. For the rest of the frame pairs, we estimate joint parameters for the revolute joint and the prismatic joint separately. For each joint type, we run RANSAC [RANSAC] for 50 iterations to select the joint parameters estimation that best fits the observation. Based on the estimated joint parameters, we can transform the feature points to expect position at time step . By computing the distance between the expected position and the observed position , we can evaluate how well this set of joint parameters describes the observation. If the distance for the revolute joint is lower, this frame pair votes for revolute, and vice versa. After estimating joint parameters for all the frame pairs, we compute the mean value for the joint axis and position for both joint types. We choose the joint type that gets the most votes as the predicted joint type for coarse prediction.
@@
Though we predict the joint type in coarse prediction, we still retain the prediction results of both joint types from coarse prediction and initialize the refinement module with those values separately. For moving vector initialization, we compute the intersection area between each part and the moving map from MonST3R [zhang2024monst3r], and then normalize it by the area of each part. We run the forward computation described in the main paper. But computing the Chamfer Distance will be very slow if the number of points is very high. Therefore, for each frame, we randomly sample 50% of points from the whole observation to compute the Chamfer Distance. The gradient will not be affected since we are optimizing the moving vector instead of the probability of being a moving part of each pixel at each frame. This strategy significantly improves the optimization speed without losing prediction accuracy. We concurrently optimize the joint prediction for both joint types for 400 iterations, using Adam optimizer [kingma2014adam] and 5e-3 learning rate.
After optimization, we have two sets of optimized joint parameters, and we need to choose the joint type. We first classify the test case as a prismatic joint if the distance between the object surface point cloud and the joint axis is larger than 0.1 meters. For the remaining cases, we compare the lowest chamfer distance of these two joint types and select the joint type that has the lower Chamfer Distance. These are the final joint type prediction results for our method. Finally, we find pixels in the moving map whose value is larger than 0.7 to be the moving points. We project these points back to 3D and find points in the object surface point cloud whose distance to the moving points is less than 0.03 meters to be the movable parts of the object. The other points in the object surface point cloud are classified as static parts of the object. We reconstruct meshes using NKSR [huang2023nksr].
@@ A.2 Baselines
Robot-See-Robot-Do (RSRD).
We follow the original pipeline of RSRD to run the experiments on our dataset. RSRD will estimate the motion of each part at each frame. Based on the part motion, we select the motion of the part that has the largest translation distance throughout the video to compute the joint parameters. For geometry reconstruction, we export the gaussian splats from the dig model in RSRD pipeline [kerr2024rsrd] and build mesh using NKSR [huang2023nksr].
Articulate Anything (AA).
We follow the original pipeline of AA to run the experiments on our dataset. AA will retrieve top-k objects from the Partnet-Mobility dataset based on the input video. Then, it selects the best one from top-k candidates [le2024articulate]. When evaluating its results, we consider them to be a failure if the correct object does not lie in the retrieved top-k candidates.
@@ Appendix B Dataset Details
We build our dataset in the SAPIEN simulator [Xiang_2020_SAPIEN]. We design a small cuboid space with a simple texture as the environment. Then, we place the object in that space and render videos while manipulating the object using the built-in controller. Fig.˜B.1 illustrates the data collection environment. The statistics are listed in Tab.˜B.1.
| Category | # Objects | Joint Type | # Videos (revolute, prismatic) | # Average Frames |
@@ C.1 Joint Estimation and Geometry Reconstruction Results on L-Dataset
We conduct a more extensive evaluation of methods, excluding RSRD, on the L-Dataset. Quantitative results on joint estimation and geometry reconstruction are shown in Tab.˜C.2 and Tab.˜C.3. Similar to the results in S-Dataset, our method outperforms AA and other ablated versions consistently on L-Dataset.
@@ C.3 Moving Map Segmentation Results
Moving map segmentation is an important intermediate results of our method. We evaluate moving map segmentation using mIOU metric. From the quantitative results in Tab.˜C.5, we find our method achieves comparable results to MonST3R [zhang2024monst3r], though our pipeline ignores the inner parts of the object, which could be moving part in the video. Without coarse prediction as initialization, our method is hard to optimize for very accurate moving map segmentation. Similarly, without automatic mask generation, our pipeline cannot optimize moving map of the video efficiently, resulting in worse moving map segmentation of the video.
@@ C.4 Camera Pose Estimation Results
| Dataset | Rot. Error(rad) | Pos. Error(m) |
| S-Dataset | 0.008 | 0.020 |
| L-Dataset | 0.011 | 0.029 |
There are already several dynamic scene reconstruction works that can infer camera poses of a video sequence. This comparison aims to demonstrate why not directly use camera estimation results from existing dynamic scene reconstruction models. We compare our camera pose prediction results with two very recent works MonST3R [zhang2024monst3r] and CUT3R [cut3r]. We transform all the camera poses to the camera coordinate of the first video frame. We compare both camera rotation error and translation error on our dataset. From the results shown in Tab.˜C.6, we can find that our method can produce more accurate camera pose estimation, particularly camera orientation, against recent works on dynamic scene understanding. We further evaluate how camera pose estimation will affect the final joint parameter estimation results on S-Dataset. In Tab.˜C.8, we find that using camera poses estimated by MonST3R will lead to a large degradation in the performance.
We also evaluate the accuracy of the strategy of aligning the first video frame to the static scan coordinate system in Tab.˜C.7. On both datasets, the camera pose estimation error is small, showing the effectiveness of our strategy of aligning the first video frame to the scan coordinate system.
@@ C.5 Moving Map Prediction Choice
In our coarse prediction module, we use MonST3R [zhang2024monst3r] to generate a moving map of the input video. A natural question is, why not use an off-the-shelf motion segmentation method? Here we compare the output of moving map prediction results from MonST3R and Segment Any Motion [huang2025segmentmotionvideos], which is a very recent work on motion segmentation on videos. In Fig.˜C.2 we can find that Segment Any Motion produces much larger error prediction results compared to MonST3R. Segment Any Motion first computes point tracks of the video, and selects moving points based on point tracks. Then, it uses these points to prompt Segment Anything 2 [ravi2024sam2] to generate the moving map of the video [huang2025segmentmotionvideos]. But in this case, the moving points lie around the edge of the door of this dishwasher. Thus, some points prompt selecting the dishwasher’s top as the moving part instead of the front door. This small position error will lead to a large error in moving map segmentation. In contrast, MonST3R computes the moving map based on optical flow and estimated camera poses. It subtracts the optical flow caused by camera motion from the final optical flow to determine which part of the video is moving. This strategy turns out to be much robust to point-track based motion segmentation methods.
@@ C.6 Discussion on Automatic Video Segmentation Strategy
Besides, we test our method using ground truth mesh-level segmentation on the input video on S-Dataset. An image is rendered from an object comprising several meshes. The mesh-level segmentation identifies different regions that are rendered from different meshes. The results in Tab.˜C.9 show that at least using automatic part segmentation is comparable to using ground truth mesh-level segmentation.
@@ C.7 Additional Real-World Results
In the main paper, we demonstrate the results of building interactable digital twins of real articulated objects with revolute joints. Here we demonstrate reconstructing articulated objects with prismatic joints in Fig.˜C.4. AA still struggles with retrieving the correct object in the Storage case, while RSRD fails to crop and cluster the object from the environment for the Drawer case.
C.8 Failure Case Analysis and Limitation
@@
From the failure modes discussed above, we believe geometry information will help to refine the segmentation results. For example, we can use recent 3D segmentation approaches like SAMPart3D [yang2024sampart3d] and PartField [partfield2025] to produce segmentation on the observed point cloud. This geometry-based segmentation can correct the errors in automatic part segmentation, which is purely image-based.